OpenStudy (anonymous):

The region bounded by y= the Gaussian function, y=0, x=0 and x=1 and is revolved about the y-axis. Find the volume of the resulting solid. anyone?

OpenStudy (anonymous):

the Gaussian function is e^-(x^2)

OpenStudy (anonymous):

\[\int\limits_{0}^{1}-e^-(x^2)\] is that the correction operation? I am confused...

OpenStudy (anonymous):

you should use "cylindrical shells". if e^-(x^2) is really the "gaussian function" the integral is : \[2\pi\int\limits\limits_{0}^{1}xe^-(x^2)dx\]